# Compactfying Electronic Wavefunctions I: Error-Mitigated Transcorrelated DMRG

### v1

@@ 1 Introduction
Efficiently addressing electronic Hamiltonians is paramount in research areas such as quantum chemistry, materials science, and condensed matter physics. A range of approaches have been employed to achieve this goal, such as full configuration interaction quantum Monte Carlo (FCIQMC) [1, 2] and various Density Matrix Renormalization Group (DMRG) formulations [3]. Given the inherent complexity of these systems, understanding electron correlation types is essential. Electron correlations are often categorized into two main types: static and dynamic.
Dynamic correlation describes the instantaneous interactions between electrons, particularly in cases where electrons occupy different spatial orbitals. This is frequently referred to as a near-degeneracy effect, significant in systems where various orbitals share similar energy levels. For example, in helium, electron correlation is predominantly dynamic, while in the $H_{2}$ molecule at the dissociation limit, correlation is entirely static, where the bonding and anti-bonding molecular orbitals become degenerate. At equilibrium in $H_{2}$, correlation primarily reflects dynamic behavior, similar to helium, yet it shifts to a static form as bond length increases. Likewise, the beryllium atom exhibits both static and dynamic correlations.
@@ 2 New Research Initiative: It from Qubit and Bootstrap for Chemistry
Inspired by string theory and quantum gravity, Wen proposed a unifying framework called string-net condensation, where collective excitations of fluctuating strings in a lattice give rise to emergent gauge fields and particles. This model explains the emergence of photons and fermions as low-energy excitations, providing a condensed matter interpretation of gauge bosons and fermions [43]. He later unified these descriptions using tensor category theory, bridging the gap between algebraic structures and physical observables [44].
@@
Inspired by this philosophy of interdisciplinary synthesis, the It from Qubit and Bootstrap for Chemistry program seeks to construct a new interdisciplinary bridge, connecting quantum information theory, conformal bootstrap, and quantum chemistry. Just as the typical theoretical physics tradition demonstrated as legendary figure such as Weinberg (sought to bring together the weak and electromagnetic forces into a single framework) and Wen unified topological field theories with condensed matter physics. Inspired by several examples in science, we aim to merge insights from high-energy physics and quantum computation to develop a systematic methodology for understanding strongly correlated systems and ab initio Hamiltonians in chemistry.
@@ 2.0.1 Program Structure
2.0.2 Transcorrelation and Its Renaissance in Quantum Chemistry
@@
The TC method’s application to strongly correlated systems—especially within the context of ab initio Hamiltonians—gained renewed attention in recent years. Advances in explicitly correlated R12/F12 methods [71, 76, 77, 78, 79, 80] and computational frameworks such as full configuration interaction quantum Monte Carlo (FCIQMC) [81, 82, 83, 84] have underscored the importance of TC techniques. Despite the challenges posed by the method’s non-Hermitian nature, its ability to improve basis-set convergence and compact the representation of wavefunctions has inspired a second wave of innovation, including its adaptation to coupled-cluster and hybrid variational algorithms [85, 86].
@@
2.0.3 Transcorrelation as a Key Element of Program Philosophy
@@ 3.2.1 The Transcorrelated Method and its Extensions
where the $\alpha_{ik}$, $\gamma_{k}$, $B_{k}$, and $A_{k}$ are non-linear parameters which need to be optimized variationally [49]. This equation is a product of two s-type Gaussian functions with a Gaussian correlation factor $\exp\left(-\gamma_{k}r^{2}\right)$. Note that the Gaussian geminals do not satisfy the cusp condition (19) but its linear combination can be approximated of the shape of the electronic wavefunction near the cusp due the dependence of the explicitly interelectronic distance. Handy successfully applied this method to small molecular systems, such as $H_{2}O$, achieving accurate results with a small number of parameters [110].
@@ 3.2.4 Recent Advances in the Transcorrelated Method
By transferring dynamic correlation from the wavefunction to the TC Hamiltonian, Luo and Alavi substantially reduced the number of walkers required in FCIQMC simulations, thereby enhancing computational efficiency. Their approach has been successfully applied across a range of systems [118, 119]. The transcorrelated method provides a versatile framework for managing the complexity of many-body quantum problems by transforming the Hamiltonian to more effectively incorporate electron correlation effects. Despite the challenges posed by its non-Hermitian nature, including complexities in computing observables beyond energy, recent developments have substantially improved its applicability. The TC method holds considerable potential, especially in the context of quantum computational algorithms, and continues to drive progress in electronic structure theory.
@@ 5 Numerical Experiments: EMTC-DMRG
As was explained in the section 3, we will test our algorithm for the Fermi-Hubbard model under some regimes:
@@
We applied our EMTC-DMRG, explained in Section 3, to the two-dimensional Fermi-Hubbard model, considering periodic boundary conditions, different lattice sizes, fillings, and interaction strengths. In our approach, as described in Section 4, we use time-independent DMRG (TI-DMRG) and map the 2D lattices to 1D using a snake-like model. We combined this with an analytical formulation of the transcorrelated transformation for the target system. Furthermore, our Hamiltonian is described in real space throughout this section, while the TC-DMRG method, with which we compare in this section, is described using a k-space Hamiltonian. The numerical experiments presented here were conducted using a Python-based code. Our Hamiltonian was implemented using the OpenFermion library, with an interface developed for the Renormalizer library, as proposed by [28]. The exact reference values used were obtained from [144, 145]. The outcomes presented on the tables, the Jastrow factor labeled as $J_{our}$ would be related to EMTC-DMRG and labeled as $J_{other}$ is related to the outcomes of the TC-DMRG presented in [29]. Our simulations were performed using an NVIDIA GeForce RTX 3050 Laptop GPU.
@@
Surprisingly, we ran some tests as shown in Table 1, where we can see that the results from a bond dimension of $m=70$ and $J=-0.15$ reach values close to the true ground state $-0.0012$. The results obtained with our method converged within $5$ to $7$ sweeps. It is important to realize that in our approach as we do not use imaginary time evolution DMRG in one of the steps or the entire procedure, we have the breaking of variational principle since we are dealing with a non-hermitian.
So, that indicates that we are reducing physical resources since a significative reduction of the value of the bond dimension. Also, in our approach, the transcorrelation makes a difference and helps to the faster convergence regarding the original model. We see that if we keep increasing the transcorrelation parameter $J$ the ground state values keep going below the true value, but in this case still is physically acceptable and we chose the value of the bond dimension closer to the exact value.
@@ 7 Conclusion and Outlook
8 Acknowledgments
BGMA expresses gratitude to Akimasa Miyake for the opportunity to spend a year in his group, for providing a challenging and independent research direction, for encouraging intellectual risks, and for his patience. We extend our sincere thanks to Jiajun Ren for his invaluable guidance in using the Renormalizer library and for the insightful discussions that significantly contributed to this work.
BGMA also acknowledges Marco A. Rodríguez-García for dedicating significant time to running numerous simulations involving the imaginary EMTC-DMRG and iTC-DMRG methods during a busy period, for making her stay at CQuIC and in Albuquerque more enjoyable, and for providing valuable feedback on this paper.
We also thank Sam McArdle and Werner Doubratz for insightful discussions. Additionally, I am grateful to Hong-Hao Tu for providing outstanding video and PDF lectures on tensor networks and for valuable discussions on projected wavefunctions.
Finally, we express our gratitude to Chung-Yun Hsieh for his careful reading of this work, thoughtful feedback, and encouraging attitude.
This work was supported by the EAGER NSF CHE-2037832 grant (September 2022 – August 2023) and CNPq grants 150963/2024-6 (BGMA) and 307626/2022-9 (AMSM).

### v2

# Compactifying Electronic Wavefunctions I: Error-Mitigated Transcorrelated DMRG

@@ Abstract
thanks: bruna.gabriellyma@gmail.comthanks: antonio.smacedo@ufpe.br
@@ 1 Introduction
Efficiently addressing electronic Hamiltonians is paramount in research areas such as quantum chemistry, materials science, and condensed matter physics. A range of approaches have been employed to achieve this goal, such as full configuration interaction quantum Monte Carlo (FCIQMC) [1, 2] and various Density Matrix Renormalization Group (DMRG) formulations [3]. Given the inherent complexity of these systems, understanding the nature of electron correlation is essential, as it dictates the accuracy and efficiency of computational methods. Since this work involves specialized quantum chemistry concepts, we provide a brief glossary of methods and key concepts in Appendix 8 to assist readers unfamiliar with these techniques.
Electron correlation is typically classified into two main types: static and dynamic, each playing a crucial role in electronic structure calculations. Dynamic correlation describes the instantaneous interactions between electrons, particularly in cases where electrons occupy different spatial orbitals. This is frequently referred to as a near-degeneracy effect, significant in systems where various orbitals share similar energy levels. For example, in helium, electron correlation is predominantly dynamic, while in the $H_{2}$ molecule at the dissociation limit, correlation is entirely static, where the bonding and anti-bonding molecular orbitals become degenerate. At equilibrium in $H_{2}$, correlation primarily reflects dynamic behavior, similar to helium, yet it shifts to a static form as bond length increases. Likewise, the beryllium atom exhibits both static and dynamic correlations.
@@
1.1 Conceptual Clarifications and Terminology
Here we briefly clarify certain terminologies used extensively in this paper to avoid potential confusion for readers unfamiliar with concepts from theoretical physics or quantum chemistry.
1.1.1 Compactification in Theoretical Physics
In theoretical physics, the term compactification typically refers to a procedure used to reduce the apparent number of dimensions of a theory, particularly common in high-energy physics, string theory, and related fields. Compactification generally involves the following steps:
  • Starting from a higher-dimensional theory.
  • Suppose that the extra dimensions are compactified, or arranged in small, typically unobservable spaces.
  • Obtaining an effective theory that appears lower-dimensional.
Thus, the compactified dimensions remain hidden from the experiments at currently accessible energy scales. In string theory, compactification plays a fundamental role, since string theories are initially formulated in higher-dimensional spacetime (10 or 11 dimensions, depending on the theory).
To reconcile this with the observable four-dimensional universe, additional spatial dimensions are compactified. Important references on compactification in string theory include [155, 156, 157, 158, 159, 160]. Common compactification manifolds in string theory are:
  • Calabi-Yau manifolds [157]: Complex shapes widely employed in compactifications from 10 dimensions down to our observed 4-dimensional spacetime.
  • Toroidal compactification (Tori) [156]: Simpler compactifications where extra dimensions form loops.
These compactifications are not merely mathematical conveniences. They critically influence the physical properties of the lower-dimensional theory, such as particle spectra, symmetries, and coupling constants.
In quantum chemistry or quantum many-body physics, the concept of compactification can also appear metaphorically in this work. In this context, compactification refers to the reduction of the complexity or effective dimensionality of electronic wavefunctions, making them more computationally tractable while retaining their essential physical characteristics.
1.1.2 Error Mitigation in Quantum Computation and EMTC-DMRG
The term error mitigation in the context of our work has a distinct meaning compared to its typical use in quantum computing. In this section, we clarify these differences explicitly.
In quantum computing, error mitigation generally refers to techniques aimed at reducing or suppressing errors in quantum computations without incurring the substantial overhead associated with full quantum error correction. Typical sources of errors include noise, decoherence, gate imperfections, or measurement inaccuracies  [122, 123]. The prominent error mitigation strategies used in quantum computing are:
  • Zero-noise extrapolation  [122, 123, 126] ;
  • Readout error mitigation  [125];
  • Quasi-probability methods  [123]
These techniques aim to enhance the reliability of quantum computations in the presence of imperfect physical implementations. In contrast, the term error mitigation as used in our paper, EMTC-DMRG, has a different conceptual context. Specifically, EMTC-DMRG explicitly focuses on mitigating intrinsic difficulties arising from the non-Hermitian nature of the transcorrelated Hamiltonian.
The transcorrelation (TC) method applies a similarity transformation to the Hamiltonian method to embed electron-electron correlations explicitly. A crucial consequence of this transformation is that the resulting Hamiltonian becomes non-Hermitian  [19, 29]. This non-Hermiticity introduces significant computational challenges.
Numerical instabilities due to non-Hermitian eigenvalue problems, especially when employing iterative algorithms such as the Davidson solver. Besides, difficulties in achieving numerical stability and reliable convergence when applying algorithms such as the DMRG.
Therefore, in EMTC-DMRG, error mitigation specifically denotes strategies devised to manage or compensate for these intrinsic algorithmic challenges rather than physical quantum hardware noise. The key methods used to achieve this goal are the following:
  • Analytical and symbolic Matrix Product Operator (MPO) constructions to minimize numerical errors. Utilization of graph-theoretic algorithms (e.g., Hopcroft-Karp) for optimal MPO construction, thus reducing computational complexity [28].
  • Using TI- DMRG with usual Davidson solver to avoid Trotter errors [127].
Additionally, optimal MPO parameters and the careful application of the Davidson solver ensure stability against numerical instabilities. In summary, EMTC-DMRG’s approach to error mitigation:
  • Does not focus on hardware-induced quantum noise.
  • Addresses numerical instabilities and approximation errors intrinsic to the non-Hermitian transformation involved.
  • Employs a combination of existing techniques to ensure accuracy and computational stability.
Thus, while the term error mitigation typically relates to quantum computational hardware issues, our work adapts this concept to the context of algorithmic and mathematical instabilities specific to the transcorrelated methodology presented herein.
@@ 2 New Research Initiative: It from Qubit and Bootstrap for Chemistry
Inspired by string theory and quantum gravity, Wen proposed a unifying framework called string-net condensation, where collective excitations of fluctuating strings in a lattice give rise to emergent gauge fields and particles. This model explains the emergence of photons and fermions as low-energy excitations, providing a condensed matter interpretation of gauge bosons and fermions [43]. He later unified these descriptions using tensor category theory, bridging the gap between algebraic structures and physical observables.
@@
Inspired by this philosophy of interdisciplinary synthesis, we propose a program called the It from Qubit and Bootstrap for Chemistry, that seeks to construct a new interdisciplinary bridge, connecting quantum information theory, conformal bootstrap, and quantum chemistry. Just as the typical theoretical physics tradition demonstrated as legendary figure such as Weinberg (sought to bring together the weak and electromagnetic forces into a single framework) and Wen unified topological field theories with condensed matter physics. Inspired by several examples in science, we aim to merge insights from high-energy physics and quantum computation to develop a systematic methodology for understanding strongly correlated systems and ab initio Hamiltonians in chemistry.
@@ 2.1 Program Structure
2.2 Transcorrelation and Its Renaissance in Quantum Chemistry
@@
The TC method’s application to strongly correlated systems—especially within the context of ab initio Hamiltonians—gained renewed attention in recent years. Advances in explicitly correlated R12/F12 methods [70, 75, 76, 77, 78, 79] and computational frameworks such as full configuration interaction quantum Monte Carlo (FCIQMC) [80, 1, 81] have underscored the importance of TC techniques. Despite the challenges posed by the method’s non-Hermitian nature, its ability to improve basis-set convergence and compact the representation of wavefunctions has inspired a second wave of innovation, including its adaptation to coupled-cluster and hybrid variational algorithms [82, 83].
@@
2.3 Transcorrelation as a Key Element of Program Philosophy
@@ 3.2.1 The Transcorrelated Method and its Extensions
where the $\alpha_{ik}$, $\gamma_{k}$, $B_{k}$, and $A_{k}$ are non-linear parameters which need to be optimized variationally [48]. This equation is a product of two s-type Gaussian functions with a Gaussian correlation factor $\exp\left(-\gamma_{k}r^{2}\right)$. Note that the Gaussian geminals do not satisfy the cusp condition (19) but its linear combination can be approximated of the shape of the electronic wavefunction near the cusp due the dependence of the explicitly interelectronic distance. Handy successfully applied this method to small molecular systems, such as $H_{2}$, achieving accurate results with a small number of parameters [106].
@@ 3.2.4 Recent Advances in the Transcorrelated Method
By transferring dynamic correlation from the wavefunction to the TC Hamiltonian, Luo and Alavi substantially reduced the number of walkers required in FCIQMC simulations, thereby enhancing computational efficiency. Their approach has been successfully applied across a range of systems [113]. The transcorrelated method provides a versatile framework for managing the complexity of many-body quantum problems by transforming the Hamiltonian to more effectively incorporate electron correlation effects. Despite the challenges posed by its non-Hermitian nature, including complexities in computing observables beyond energy, recent developments have substantially improved its applicability. The TC method holds considerable potential, especially in the context of quantum computational algorithms, and continues to drive progress in electronic structure theory.
@@ 5 Numerical Experiments: EMTC-DMRG
The numerical validation presented here focuses on the transcorrelated Fermi-Hubbard model, which serves as a rigorous testbed for our approach. Given that EMTC-DMRG is designed to handle transcorrelated fermionic systems, further case studies would provide similar results, as the key improvements—error mitigation and computational efficiency—are independent of the specific model choice. As was explained in the section 4, we will test our algorithm for the Fermi-Hubbard model under some regimes:
@@
We applied our EMTC-DMRG, explained in Section 3, to the two-dimensional Fermi-Hubbard model, considering periodic boundary conditions, different lattice sizes, fillings, and interaction strengths. In our approach, as described in Section 4, we use time-independent DMRG (TI-DMRG) and map the 2D lattices to 1D using a snake-like model. We combined this with an analytical formulation of the transcorrelated transformation for the target system. Furthermore, our Hamiltonian is described in real space throughout this section, while the TC-DMRG method, with which we compare in this section, is described using a k-space Hamiltonian.
The numerical experiments presented here were conducted using a Python-based code. Our Hamiltonian was implemented using the OpenFermion library, with an interface developed for the Renormalizer library, as proposed by [28]. The exact reference values used were obtained from [143, 144]. The outcomes presented on the tables, the Jastrow factor labeled as $J_{our}$ would be related to EMTC-DMRG and labeled as $J_{other}$ is related to the outcomes of the TC-DMRG presented in [29]. Our simulations were performed using an NVIDIA GeForce RTX 3050 Laptop GPU.
@@
Surprisingly, we ran some tests as shown in Table 1, where we can see that the results from a bond dimension of $m=70$ and $J=-0.15$ reach values close to the true ground state $-0.0012$. The results obtained with our method converged within $5$ to $7$ sweeps. It is important to realize that in our approach, as we do not use imaginary time evolution DMRG in one of the steps or the entire procedure, we have the breaking of the variational principle since we are dealing with a non-hermitian.
So, that indicates that we are reducing physical resources since a significant reduction of the value of the bond dimension. Also, in our approach, the transcorrelation makes a difference and helps to the faster convergence regarding the original model. We see that if we keep increasing the transcorrelation parameter $J$, the ground state values keep going below the true value, but in this case still is physically acceptable, and we chose the value of the bond dimension closer to the exact value.
To enhance readability, we provide visual representations of these numerical results in Figures 2,3, and 4 which illustrate the trends observed in Tables 1,2, and 3.
@@
Figure 2: Ground state energy vs. bond dimension for the $3\times 3$ Fermi-Hubbard model, corresponding to Table 1.
@@
Figure 3: Ground state energy vs. bond dimension for the $4\times 4$ Fermi-Hubbard model, corresponding to Table 2.
@@
Figure 4: Ground state energy vs. bond dimension for the $4\times 4$ Fermi-Hubbard model at $U/t=4$, corresponding to Table 3.
@@
To summarize the differences between our EMTC-DMRG and existing transcorrelated DMRG approaches, we provide a comparative in Table 7. This table highlights the key methodological distinctions, computational trade-offs, and advantages of each approach. Notably, our method improves numerical stability through symbolic optimization and achieves faster convergence without requiring imaginary-time evolution, unlike previous TC-DMRG formulations.
Algorithm EMTC-DMRG TC-DMRG [29] TC-DMRG [32]
Main Approach Symbolic optimization; TI-DMRG applied to a TC Hamiltonian described in real space to mitigated errors. Uses TI-DMRG and iTD- DMRG and applied to TC Hamiltonian described in k-space. Time-independent DMRG with a generalized Davidson solver for non-Hermitian Hamiltonians
Key Features Optimal exact MPO; Analytical TC-FH Hamiltonian; Standard Davidson solver Exact MPO; Analytical TC- FH Hamiltonian; Standard Davidson solver. Compressed MPO; Approximated TC Hamiltonian formulation; Generalized Davison solver.
Challenges Addressed Reducing long-range interactions; Show that TC helps to improve the 2D DMRG performance to compute ground state and in some cases outperforms [29]. Reducing long-range interactions; Show that TC helps to improve the 2D DMRG performance to compute ground state. Shows that they TC formulation improves the efficiency and accuracy in computingt ground/excited state of some molecules.
Target Systems 2D Fermi-Hubbard model with pbc. 2D Fermi-Hubbard model with pbc. Molecular systems.
Computational Efficiency Significant reduction of the resources like bond dimension $m$ in relation to [29]. Significant reduction of the resources like bond dimension $m$ in relation to TI-DMRG. Accelerated basis set convergence for molecular systems.
Variational Principle Does not apply here. Applies here due the use of imaginary time evolution. Applies here due the use of Generalized Davidson solver for non-Hermitian Hamiltonian.
Limitations Not optimal for very large lattices without a k-space Hamiltonian Requires imaginary-time evolution, and k-space formulation of the Hamiltonian even for modest size lattices. Requires Generalized Davidson solver; Only suitable for molecular systems.
Table 7: Comparison of Transcorrelated DMRG Methods.
Unlike previous TC-DMRG formulations, our method eliminates the need for imaginary-time evolution, reducing computational overhead while achieving comparable accuracy. This is reflected in Table 7.
@@ 7 Conclusion and Outlook
Acknowledgments
BGMA expresses gratitude to Akimasa Miyake for the opportunity to spend a year in his group, for providing a challenging and independent research direction, for encouraging intellectual risk-taking, and for his patience.
We extend our sincere thanks to Jiajun Ren for his invaluable guidance in using the Renormalizer library and for the insightful discussions that significantly contributed to this work. BGMA also acknowledges Marco A. Rodríguez-García for dedicating substantial time to running numerous simulations involving the iEMTC-DMRG and iTC-DMRG methods and for making her stay at CQuIC and in Albuquerque more enjoyable.
The authors thank Aron W. Cummings for his excellent feedback and constructive criticism, which improved the quality of this manuscript.
We also thank Sam McArdle and Werner Doubratz for their discussions. Additionally, we appreciate Hong-Hao Tu for providing excellent video and PDF lectures on tensor networks, as well as for valuable discussions on projected wavefunctions.
We thank Guido Raos, Piero Macchi, Mosè Casalegno, and Alessandro Genoni for their guidance on electron correlation problems. Finally, we express our gratitude to Chung-Yun Hsieh for carefully reviewing parts of this work, providing thoughtful feedback, and offering encouraging support.
This work was supported by the EAGER NSF CHE-2037832 grant (BGMA), CNPq grants 150963/2024-6 (BGMA), and 307626/2022-9 (AMSM).